OpenStudy (anonymous):

The sum of the two digits of a number is 16. The number formed by reversing the digits is 18 more than the original number. Determine the original number. Let t = the tens digit, u = the units digit, and u + t = 16. Which of the following equations would complete the system? 9t - 9u = 18 9u - 9t = 18 tu = ut + 18

OpenStudy (anonymous):

@sourwing?

OpenStudy (anonymous):

10t + u = original number when reversed: 10u + t = new number new number = 18 + oringal number, so 10u + t = 18 + 10t + u

OpenStudy (anonymous):

that's is for the first part. You try the second part

OpenStudy (anonymous):

how did you get 10?

OpenStudy (anonymous):

say you have 98. wouldn't 98 = 10(9) + 8? because the 9 is the tens digit

OpenStudy (anonymous):

oh wait i get lol sorry

OpenStudy (anonymous):

so isnt it the middle one

OpenStudy (anonymous):

@mathmale will you help

OpenStudy (mathmale):

I'll try! But it'll be a minute or two before I can get back to you.

OpenStudy (anonymous):

ok thanks!

OpenStudy (mathmale):

I haven't read all of the discussion you've had just recently. Would you mind telling me as accurately as you can what you need to know to solve this problem, and how I may be able to help with this?

OpenStudy (mathmale):

If t=10's digit and u=1's digit, then tu represents the original number. It just so happens that t+u=16. Just supposing that t=7 and u=9, then t+u=16, right? tu would be 79 and ut would be 97. Give this some thought. Actually, I've just made educated guesses and believe that was enough to give me the answer: ut=the original number = 79. Still, you must be able to write equations that lead you to the correct values of u and t. I'm really sorry, but I need to get off my computer pronto. I hope these leads are sufficient to help you come up with proper equations for the identification of t and u.